OpenStudy (anonymous):

To calculate the money spent on gas, use the average price of gas $3.50 per gallon, and the average distance per gallon in 25 miles. Let's assume that the average fuel load for a typical car is 12 gallons per tank. How many times would you need to fuel-up on your trip?

OpenStudy (heisenberg):

To answer this question, you need to know how many miles your trip is.

OpenStudy (anonymous):

The trip is 40hours and 51 minutes

OpenStudy (heisenberg):

ok, well now that brings up another question! lol. how fast are we traveling? if you just said we were traveling '25 miles' that would be easy, but instead you gave me a time which could be any amount of miles.

OpenStudy (heisenberg):

focus on the dimensions (the canceling of units) and you will see how this makes sense: \[\frac{miles}{gallon} * \frac{gallons}{tank} = \frac{miles}{tank}\]

OpenStudy (heisenberg):

divide your total trip distance (which is in miles) by this answer. you can see the dimensions in our favour: \[miles / \frac{miles}{tank} = tanks\] this is the number of tanks required for our trip. the initial money per gallon info is useless

OpenStudy (anonymous):

I just wante to let you know that im in the 8th grade. And im not familiar with this.

OpenStudy (heisenberg):

that's perfectly fine. just look at the units (miles, gallons, etc) like fractions and you can see how the numerator cancels with the denominator

OpenStudy (anonymous):

clearly there is not enough information to solve this

OpenStudy (anonymous):

in the future, type out the entire question. even part 1) part 2) part 3)....